OpenStudy (anonymous):

Given f(x) = (x-1)/2, solve for f^-1 (5)

OpenStudy (funnyguy2):

is this a test

OpenStudy (anonymous):

@funnyguy2 classwork, I'm just really confused

OpenStudy (anonymous):

Hi welcome to openstudy!! FIRST, set f(x) = y so your equation will be \(\sf y=\frac{ (x-1)}{2}\) SECOND, switch the places of x and y since you are looking for the inverse of the function: you'll have: \(\sf x=\frac{ (y-1)}{2}\) THIRD, solve the equation in terms of y, can you do this for me ? :)

OpenStudy (anonymous):

@Data_LG2 I am a little confused what you mean by that. do you want me to multiply -1 to both sides? So then Ill have x= y/2?

OpenStudy (anonymous):

nope, try arranging the equation so that y will be on the left side instead of x

OpenStudy (anonymous):

for example: 

OpenStudy (anonymous):

So I will multiply y to both sides which will give me yx=-1/2 and then I will divide by x?

OpenStudy (anonymous):